OpenStudy (anonymous):

An angle is 6 degrees more than twice its complement. What is the measure of the angle?

OpenStudy (anonymous):

Let the measure of one angle be x degrees therefore measure of its complement would be (x+ 6) degrees Since the sum of complementary angle = 90 degrees therefore x+ (x+6) = 90 2x+6 = 90 2x = 90-6 2x= 84 x= 84/2 x= 42 therefore x+ 6 = 42 + 6 = 48 hence the measure of the angles are 42 degrees and 48 degrees.